OpenStudy (anonymous):

Alex has twice as much money as Jennifer. Jennifer has $6 less than Shannon. Together they have $54. How much money does Alex have?

OpenStudy (anonymous):

i am terrible at these sort of puzzles

OpenStudy (anonymous):

Money Alex has = A Money Jennifer has = J Money Shannon has = S A = 2J Equation (1) S - J = 6 Equation (2) A + S + J = 54 Equation (3) Solve the 3 equations.

OpenStudy (anonymous):

54 divided by 3= x x - 6 equals what jennifer has x-6 times 2 = what alex has add them all together to check your answer

OpenStudy (anonymous):

j=12 a=24 2=18

OpenStudy (anonymous):

Yup. :) That's right.